OpenStudy (anonymous):

what number best completes this sequence 3 94 5 25 20 21 441

OpenStudy (anonymous):

Pattern: Square the term, then take 5, add one, and square it again, and so on. So 3x3=9, 9-5=4, 4+1=5, and 5x5=25. Next number= 441-5=436